# mFabric: An Efficient and Scalable Fabric for Mixture-of-Experts Training

### v2

@@ Abstract.
Mixture-of-Expert (MoE) models outperform conventional models by selectively activating different subnets, named experts, on a per-token basis. This gated computation generates dynamic communications that cannot be determined beforehand, challenging the existing GPU interconnects that remain static during distributed training. In this paper, we advocate for a first-of-its-kind system, called mFabric, that unlocks topology reconfiguration during distributed MoE training. Towards this vision, we first perform a production measurement study and show that the MoE dynamic communication pattern has strong locality, alleviating the need for global reconfiguration. Based on this, we design and implement a regionally reconfigurable high-bandwidth domain that augments existing electrical interconnects using optical circuit switching (OCS), achieving scalability while maintaining rapid adaptability. We build a fully functional mFabric prototype with commodity hardware and a customized collective communication runtime. Our prototype trains state-of-the-art MoE models with in-training topology reconfiguration across 32 A100 GPUs. Large-scale packet-level simulations show that mFabric achieves performance comparable to a non-blocking fat-tree fabric while boosting the networking cost efficiency (e.g., performance per dollar) of four representative MoE models by 1.2$\times$–1.5$\times$ and 1.9$\times$–2.3$\times$ at 100 Gbps and 400 Gbps link bandwidths, respectively.
*Corresponding authors: zhizhenz@mit.edu (Z. Zhong), kaichen@cse.ust.hk (K. Chen)
submissionid: #1conference: SIGCOMM’25; September 8 - 11, 2025; Coimbra, Portugal
@@ 1. Introduction
Based on this insight, we introduce mFabric, a novel system designed to overcome these challenges by enabling efficient topology reconfiguration during distributed MoE training. At the core of mFabric is a regionally reconfigurable high-bandwidth domain based on millisecond-scale reconfigurable OCS that sits at the boundary between scale-up and scale-out networks. This design augments the existing static electrical interconnect with the capability of rapid regional reconfiguration while still preserving its scalability (§4).
mFabric contains the following key components: 1) mFabric exploits the partially predictable nature of all-to-all communications to track regional traffic demands at runtime (§5.1); 2) Based on the obtained demands, mFabric uses a greedy algorithm that generates the tailored network topology and reconfigures the OCS to realize it (§5.2); 3) With the reconfigured topology, mFabric uses a customized collective communication runtime to orchestrate inter-host DP and EP communications on EPS and regional OCS fabrics (§5.3).
To demonstrate mFabric, we build a fully functional prototype with 32 Nvidia A100 GPUs, 16 Mellanox NICs (100, [n. d.]b), a Polatis millisecond OCS (pol, [n. d.]), and an Ethernet switch. In addition, we develop a custom collective communication runtime based on Nvidia collective communications library (NCCL) (ncc, [n. d.]) to support in-training topology reconfigurations. Using this prototype, we successfully demonstrated the benefits of mFabric on state-of-the-art MoE models (§6).
To evaluate the performance of mFabric at scale, we perform packet-level simulations using four representative real-world MoE models. Our results reveal that mFabric outperforms the state-of-the-art fabrics, exhibiting a training speed comparable to Rail-optimized (rai, [n. d.]) and Fat-tree (Al-Fares et al., 2008) fabrics while significantly improving cost-efficiency. Specifically, mFabric improves networking infrastructure cost-efficiency by 1.2$\times$–1.5$\times$ (1.9$\times$–2.3$\times$) compared to Fat-tree and 1.4$\times$–1.5$\times$ (2.3$\times$–2.4$\times$) compared to Rail-optimized for 100Gbps (400Gbps) links. We also observe that mFabric outperforms TopoOpt (Wang et al., 2023) by up to 2.5$\times$ and supports scalability with the cluster size increasing to 30K+ GPUs (§7). When connected to co-packaged optical ports attached to GPUs, mFabric augments high-radix scale-up systems like NVL72 by 1.3$\times$8).
As the first system to enable in-training topology reconfiguration for distributed MoE training, we will release our code by the time of publication.
@@ 2.1. Distributed MoE Training
Expert Parallelism (EP). In MoE models, different experts in an MoE block are allocated to different GPUs (Fedus et al., 2022; Lepikhin et al., 2020; Liu et al., 2023b) (Figure 1(b)). Since each GPU needs to send its local states to other experts and receive remote states from other GPUs, the dispatching of intermediate hidden states and the collection of expert outputs are performed via two all-to-all communications. EP’s all-to-all communication is non-uniform and non-deterministic across different training iterations (§3).
Traffic volume of different parallelisms. We use Megatron-LM (Shoeybi et al., 2020) to profile three state-of-the-art MoE models (Mixtral 8$\times$7B MoE (mix, [n. d.]c), Llama-MoE (Zhu et al., 2024) and Qwen-MoE (qwe, [n. d.])) [^1] The detailed model configurations are shown in Table 3 in Appendix §A. and measure the total amount of data transfer. We plot the distribution of traffic volume in one MoE training iteration in Figure 2. For Mixtral 8$\times$7B, we observe that TP generates the highest traffic volume (60% of the total volume). The EP generates the second-highest traffic volume (30%), leaving PP and DP contributing less than 6%. For LLaMA-MoE and Qwen-MoE, we find that EP becomes the most communication intensive (more than 80%). This is because the size of the largest layer (i.e., expert) fits into a single GPU memory.
2.2. GPU Interconnects
@@ 3. Production Measurements
Unlike conventional parallelism (e.g., TP, PP, and DP) with deterministic communication patterns, EP’s communications are determined by the gate unit at runtime due to the semantic heterogeneity of the input tokens. To understand the dynamics of EP traffic patterns, we profile Mixtral 8$\times$7B (mix, [n. d.]c) in a production data center [^2] For detailed hardware configurations, please refer to Appendix §A., using a hybrid parallelism that combines an EP degree of 8, TP degree of 4, PP degree of 4 at a sequence length of 4096, and micro-batch size of 8 (Dubey et al., 2024).
@@
4. mFabric Architecture Design
So far, we have shown that MoE introduces unique traffic patterns that are temporally non-deterministic and spatially non-uniform. Now, an important question is How to design a network architecture that best serves the requirements of distributed MoE training? In this section, we first study an ideal yet practical fabric for distributed MoE training (§4.1). Then, we present our key proposal in mFabric, the regionally reconfigurable high-bandwidth domain with OCS (§4.2).
@@ 4.1. Towards An Ideal yet Practical Fabric
Figure 6. The mFabric network architecture.
Landing the ideal fabric in practice. To reconcile these challenges, we leverage the key observation that despite the non-deterministic and non-uniform characteristics of MoE all-to-all traffic, there is a strong locality for traffic variations as MoE blocks are normally placed in a pipeline (Shoeybi et al., 2020; dee, [n. d.]). Therefore, instead of building a globally reconfigurable OCS fabric, we propose designing several regionally reconfigurable OCS networks (described next in §4.2). Figure 6 depicts the network architecture of mFabric. By partitioning the network into several domains where the communication locality is strong, each regional OCS rapidly adjusts to traffic demands of EP without the complexity of global reconfiguration. By implementing regionally reconfigurable OCS networks, we leverage this locality to achieve fast reconfiguration within smaller, manageable regions. This approach balances the need for reconfigurability with practical hardware limitations, effectively supporting the dynamic communication patterns of MoE training.
@@ 4.2. Regionally Reconfigurable OCS
mFabric, as the first fabric to support the in-training topology reconfiguration, highlights the core idea of building regionally reconfigurable OCS to offload dynamic EP traffic in the existing electrical fabric. By leveraging the strong locality inherent in MoE’s all-to-all traffic, we partition the network into regions where communication demands are non-deterministic among expert layers. This regional approach allows for rapid reconfiguration within each partition, effectively overcoming the fundamental trade-off between reconfiguration speed and port count in OCS technology. By focusing on regional reconfigurability, mFabric achieves scalability while maintaining rapid adaptability to dynamic communication patterns of MoE training, alleviating the complexity of global network reconfiguration.
Where to deploy regionally reconfigurable OCS? To best serve the regional sparse all-to-all traffic patterns featured in distributed MoE training, a regional OCS is connected to a cluster of GPU servers where each server splits its NICs between EPS and OCS [^7] In the near term, mFabric’s regional OCS leverages the optical transceivers through the NIC for deployment readiness. In the long term, as the co-package optics (CPO) becomes widely adopted, mFabric’s regional OCS is compatible with the commodity optical I/O solutions (e.g., TeraPHY from Ayar Labs (aya, [n. d.])) that is directly connected to the computing chip (e.g., GPU, TPU, etc.). We leave more discussions in Appendix E.1.. Given the fact that today’s fast OCSes with millisecond-scale reconfiguration delay support up to 500 ports (Table 1) as well as a typical server contains eight NICs, a reconfigurable high-bandwidth domain interconnect through an OCS supports around 80 to 250 servers (each server assigning two to six NICs to OCS).
@@
How to reconfigure the topology? In mFabric, the OCS fabric is arranged into multiple isolated slices for each reconfigurable high-bandwidth region. Therefore, the regional OCS topology is controlled by its localized topology controller, which frequently collects traffic demands from the host servers. Within each training iteration are four all-to-all communications with the same or transposed traffic pattern. However, these traffic patterns are non-deterministic across different training iterations. Hence, we need to develop a mechanism to reconfigure the topology tailored to traffic patterns. The regional topology reconfiguration means that mFabric does not require a centralized topology controller, which avoids the scalability concerns of the control plane.
Towards a mixed optical-electrical fabric. mFabric seeks the best fit between the traffic patterns of the parallelization strategies and the corresponding switching technologies. Therefore, mFabric uses server-scale NVSwitch for tensor parallelism, regionally reconfigurable OCS for expert parallelism, and large-scale EPS for data parallelism and pipeline parallelism. To distribute data movement tasks on different fabrics, a new collective communication library that supports topology reconfiguration is needed.
5. mFabric System Implementation
Figure 7. mFabric system implementation.
To enable the aforementioned mFabric architecture, we design and implement several core components of the control and data planes. As shown in Figure 7, mFabric’s implementation contains a traffic monitor that keeps track of the traffic demands in EP to characterize subsequent all-to-all communications (§5.1). Based on the monitored traffic demands, multiple decentralized topology controllers generate and enforce topologies for regional OCSes (§5.2). Finally, a collective communication manager is responsible to steer the traffic in the mFabric fabric (§5.3).
@@ 5.1. All-to-All Traffic Characterization
In each MoE block, there are four all-to-all communication phases during each training iteration: two in the forward pass and two in the backward pass (Figure 1(b)). The first all-to-all communication occurs after the gate unit computation. The output from the gate unit999 The actual output of the gate unit is the dispatching probability distribution for each token. The expert load is derived directly from the probability distribution using the top-k parameter. across the EP groups determines the traffic matrix for this communication phase. These four all-to-all traffic matrices are strictly the same or transposed due to the symmetry of the token dispatching and collection process. For communication in the first all-to-all in FP, given that the mFabric network architecture is provisioned with millisecond-scale reconfigurable OCS, there are two options. mFabric reconfigures the OCS while blocking the training process, as this reconfiguration time cannot be hidden in the computation. On the other hand, if mFabric opts to utilize a random topology or reuse topology from previous MoE layers, it cannot benefit from the efficient circuits schedule.
Besides, we observe the partial predicabilty of the FP’s first all-to-all, which offers an opportunity to proactively reconfigure the OCS for it in advance (e.g., in attention computation phase). We offload the details of the prediction algorithm for all-to-all traffic demand in §B.1. Note that mFabric does not introduce extra demand monitoring overhead as the state-of-the-art MoE training framework already contains a mechanism to collect this information (meg, [n. d.]) to perform on-demand all-to-all transmission.
5.2. Topology Reconfiguration
In mFabric, finding an optimal topology and deriving an optical schedule is an NP-hard problem (Foerster et al., 2018). We address this challenge by introducing a lightweight greedy algorithm. The key insight is that all-to-all communication time is determined by the delay of largest transfers, which implies that the corresponding GPU pair should be allocated with more circuits. Thus, we identify the communication pairs with the longest transmission time in each iteration and assign them with direct optical links in the OCS topology. The detailed OCS reconfiguration algorithm is shown in Algorithm 1.

### v3

# MixNet: A Runtime Reconfigurable Optical-Electrical Fabric for Distributed Mixture-of-Experts Training

@@ Abstract.
Mixture-of-Expert (MoE) models outperform conventional models by selectively activating different subnets, named experts, on a per-token basis. This gated computation generates dynamic communications that cannot be determined beforehand, challenging the existing GPU interconnects that remain static during distributed training. In this paper, we advocate for a first-of-its-kind system, called MixNet, that unlocks topology reconfiguration during distributed MoE training. Towards this vision, we first perform a production measurement study and show that the MoE dynamic communication pattern has strong locality, alleviating the need for global reconfiguration. Based on this, we design and implement a regionally reconfigurable high-bandwidth domain that augments existing electrical interconnects using optical circuit switching (OCS), achieving scalability while maintaining rapid adaptability. We build a fully functional MixNet prototype with commodity hardware and a customized collective communication runtime. Our prototype trains state-of-the-art MoE models with in-training topology reconfiguration across 32 A100 GPUs. Large-scale packet-level simulations show that MixNet achieves performance comparable to a non-blocking fat-tree fabric while boosting the networking cost efficiency (e.g., performance per dollar) of four representative MoE models by 1.2$\times$–1.5$\times$ and 1.9$\times$–2.3$\times$ at 100 Gbps and 400 Gbps link bandwidths, respectively.
Network Architecture, Mixture-of-Experts, Optical Circuit Switching, AI Infrastructure
journalyear: 2025copyright: acmlicensedconference: ACM SIGCOMM 2025 Conference; September 8–11, 2025; Coimbra, Portugalbooktitle: ACM SIGCOMM 2025 Conference (SIGCOMM ’25), September 8–11, 2025, Coimbra, Portugaldoi: 10.1145/3718958.3750465isbn: 979-8-4007-1524-2/25/09ccs: Networks Network architecturesccs: Networks Data center networks
@@ 1. Introduction
Based on this insight, we introduce MixNet, a novel system designed to overcome these challenges by enabling efficient topology reconfiguration during distributed MoE training. At the core of MixNet is a regionally reconfigurable high-bandwidth domain based on millisecond-scale reconfigurable OCS that sits at the boundary between scale-up and scale-out networks. This design augments the existing static electrical interconnect with the capability of rapid regional reconfiguration while still preserving its scalability (§4).
MixNet contains the following key components: 1) MixNet exploits the partially predictable nature of all-to-all communications to track regional traffic demands at runtime (§5.1); 2) Based on the obtained demands, MixNet uses a greedy algorithm that generates the tailored network topology and reconfigures the OCS to realize it (§5.2); 3) With the reconfigured topology, MixNet uses a customized collective communication runtime to orchestrate inter-host DP and EP communications on EPS and regional OCS fabrics (§5.3).
To demonstrate MixNet, we build a fully functional prototype with 32 Nvidia A100 GPUs, 16 Mellanox NICs (100, [n. d.]b), a Polatis millisecond OCS (pol, [n. d.]), and an Ethernet switch. In addition, we develop a custom collective communication runtime based on Nvidia collective communications library (NCCL) (ncc, [n. d.]) to support in-training topology reconfigurations. Using this prototype, we successfully demonstrated the benefits of MixNet on state-of-the-art MoE models (§6).
To evaluate the performance of MixNet at scale, we perform packet-level simulations using four representative real-world MoE models. Our results reveal that MixNet outperforms the state-of-the-art fabrics, exhibiting a training speed comparable to Rail-optimized (rai, [n. d.]) and Fat-tree (Al-Fares et al., 2008) fabrics while significantly improving cost-efficiency. Specifically, MixNet improves networking infrastructure cost-efficiency by 1.2$\times$–1.5$\times$ (1.9$\times$–2.3$\times$) compared to Fat-tree and 1.4$\times$–1.5$\times$ (2.3$\times$–2.4$\times$) compared to Rail-optimized for 100Gbps (400Gbps) links. We also observe that MixNet outperforms TopoOpt (Wang et al., 2023) by up to 2.5$\times$ and supports scalability with the cluster size increasing to 30K+ GPUs (§7). When connected to co-packaged optical ports attached to GPUs, MixNet augments high-radix scale-up systems like NVL72 by 1.3$\times$8).
For more information, please visit the website: https://mixnet-project.github.io/.
@@ 2.1. Distributed MoE Training
Models Mixtral LLaMA-MoE Qwen-MoE
Size 8$\times$7B 6.7B 14.3B
# of MoE blocks 32 32 24
# experts 8 16 64
EP degree 8 16 16
TP degree 4 1 1
PP degree 4 4 4
Seq. len. 4096 4096 4096
Micro-batch size 8 8 8
Table 1. State-of-the-art MoE training configurations.
@@
Expert Parallelism (EP). In MoE models, different experts in an MoE block are allocated to different GPUs (Fedus et al., 2022; Lepikhin et al., 2020; Liu et al., 2023b; Li et al., 2024) (Figure 1(b)). Since each GPU needs to send its local states to other experts and receive remote states from other GPUs, the dispatching of intermediate hidden states and the collection of expert outputs are performed via two all-to-all communications. EP’s all-to-all communication is non-uniform and non-deterministic across different training iterations (§3).
@@
Traffic volume of different parallelisms. We use Megatron-LM (Shoeybi et al., 2020) to profile three state-of-the-art MoE models (Mixtral 8$\times$7B MoE (mix, [n. d.]c), Llama-MoE (Zhu et al., 2024) and Qwen-MoE (qwe, [n. d.])) and measure the total amount of data transfer. The detailed model configurations are shown in Table 1. We plot the distribution of traffic volume in one MoE training iteration in Figure 2. For Mixtral 8$\times$7B, we observe that TP generates the highest traffic volume (60% of the total volume). The EP generates the second-highest traffic volume (30%), leaving PP and DP contributing less than 6%. For LLaMA-MoE and Qwen-MoE, we find that EP becomes the most communication intensive (more than 80%). This is because the size of the largest layer (i.e., expert) fits into a single GPU memory.
2.2. GPU Interconnects
@@ 3. Production Measurements
Unlike conventional parallelism (e.g., TP, PP, and DP) with deterministic communication patterns, EP’s communications are determined by the gate unit at runtime due to the semantic heterogeneity of the input tokens. To understand the dynamics of EP traffic patterns, we profile Mixtral 8$\times$7B (mix, [n. d.]c) in a production data center, using a hybrid parallelism that combines an EP degree of 8, TP degree of 4, PP degree of 4 at a sequence length of 4096, and micro-batch size of 8 (Dubey et al., 2024).
The production fabric. We use a Certified Nvidia DGX SuperPOD platform (nvi, [n. d.]b) with 128 H800 GPUs and 128 ConnectX-7 400Gbps NICs in a production data center. The compute fabric is connected in a rail-optimized topology (rai, [n. d.]). We use NCCL (ncc, [n. d.]) to optimize communications on the DGX platform.
@@
4. MixNet Architecture Design
So far, we have shown that MoE introduces unique traffic patterns that are temporally non-deterministic and spatially non-uniform. Now, an important question is How to design a network architecture that best serves the requirements of distributed MoE training? In this section, we first study an ideal yet practical fabric for distributed MoE training (§4.1). Then, we present our key proposal in MixNet, the regionally reconfigurable high-bandwidth domain with OCS (§4.2).
@@ 4.1. Towards An Ideal yet Practical Fabric
Figure 6. The MixNet network architecture.
Landing the ideal fabric in practice. To reconcile these challenges, we leverage the key observation that despite the non-deterministic and non-uniform characteristics of MoE all-to-all traffic, there is a strong locality for traffic variations as MoE blocks are normally placed in a pipeline (Shoeybi et al., 2020; dee, [n. d.]). Therefore, instead of building a globally reconfigurable OCS fabric, we propose designing several regionally reconfigurable OCS networks (described next in §4.2). Figure 6 depicts the network architecture of MixNet. By partitioning the network into several domains where the communication locality is strong, each regional OCS rapidly adjusts to traffic demands of EP without the complexity of global reconfiguration. By implementing regionally reconfigurable OCS networks, we leverage this locality to achieve fast reconfiguration within smaller, manageable regions. This approach balances the need for reconfigurability with practical hardware limitations, effectively supporting the dynamic communication patterns of MoE training.
@@ 4.2. Regionally Reconfigurable OCS
MixNet, as the first fabric to support the in-training topology reconfiguration, highlights the core idea of building regionally reconfigurable OCS to offload dynamic EP traffic in the existing electrical fabric. By leveraging the strong locality inherent in MoE’s all-to-all traffic, we partition the network into regions where communication demands are non-deterministic among expert layers. This regional approach allows for rapid reconfiguration within each partition, effectively overcoming the fundamental trade-off between reconfiguration speed and port count in OCS technology. By focusing on regional reconfigurability, MixNet achieves scalability while maintaining rapid adaptability to dynamic communication patterns of MoE training, alleviating the complexity of global network reconfiguration.
Where to deploy regionally reconfigurable OCS? To best serve the regional sparse all-to-all traffic patterns featured in distributed MoE training, a regional OCS is connected to a cluster of GPU servers where each server splits its NICs between EPS and OCS [^5] In the near term, MixNet’s regional OCS leverages the optical transceivers through the NIC for deployment readiness. In the long term, as the co-package optics (CPO) becomes widely adopted, MixNet’s regional OCS is compatible with the commodity optical I/O solutions (e.g., TeraPHY from Ayar Labs (aya, [n. d.])) that is directly connected to the computing chip (e.g., GPU, TPU, etc.).. Given the fact that today’s fast OCSes with millisecond-scale reconfiguration delay support up to 500 ports (Table 2) as well as a typical server contains eight NICs, a reconfigurable high-bandwidth domain interconnect through an OCS supports around 80 to 250 servers (each server assigning two to six NICs to OCS).
@@
How to reconfigure the topology? In MixNet, the OCS fabric is arranged into multiple isolated slices for each reconfigurable high-bandwidth region. Therefore, the regional OCS topology is controlled by its localized topology controller, which frequently collects traffic demands from the host servers. Within each training iteration are four all-to-all communications with the same or transposed traffic pattern. However, these traffic patterns are non-deterministic across different training iterations. Hence, we need to develop a mechanism to reconfigure the topology tailored to traffic patterns. The regional topology reconfiguration means that MixNet does not require a centralized topology controller, which avoids the scalability concerns of the control plane.
Towards a mixed optical-electrical fabric. MixNet seeks the best fit between the traffic patterns of the parallelization strategies and the corresponding switching technologies. Therefore, MixNet uses server-scale NVSwitch for tensor parallelism, regionally reconfigurable OCS for expert parallelism, and large-scale EPS for data parallelism and pipeline parallelism. To distribute data movement tasks on different fabrics, a new collective communication library that supports topology reconfiguration is needed.
5. MixNet System Implementation
Figure 7. MixNet system implementation.
To enable the aforementioned MixNet architecture, we design and implement several core components of the control and data planes. As shown in Figure 7, MixNet’s implementation contains a traffic monitor that keeps track of the traffic demands in EP to characterize subsequent all-to-all communications (§5.1). Based on the monitored traffic demands, multiple decentralized topology controllers generate and enforce topologies for regional OCSes (§5.2). Then, a collective communication manager is responsible to steer the traffic in the MixNet fabric (§5.3). In addition, we discuss the failure handling mechanism in (§5.4).
@@ 5.1. All-to-All Traffic Characterization
In each MoE block, there are four all-to-all communication phases during each training iteration: two in the forward pass and two in the backward pass (Figure 1(b)). The first all-to-all communication occurs after the gate unit computation. The output from the gate unit777 The actual output of the gate unit is the dispatching probability distribution for each token. The expert load is derived directly from the probability distribution using the top-k parameter. across the EP groups determines the traffic matrix for this communication phase. These four all-to-all traffic matrices are strictly the same or transposed due to the symmetry of the token dispatching and collection process. For communication in the first all-to-all in FP, given that the MixNet network architecture is provisioned with millisecond-scale reconfigurable OCS, there are two options. MixNet reconfigures the OCS while blocking the training process, as this reconfiguration time cannot be hidden in the computation. On the other hand, if MixNet opts to utilize a random topology or reuse topology from previous MoE layers, it cannot benefit from the efficient circuits schedule.
Besides, we observe the partial predicabilty of the FP’s first all-to-all, which offers an opportunity to proactively reconfigure the OCS for it in advance (e.g., in attention computation phase). We offload the details of the prediction algorithm for all-to-all traffic demand in §B.1. Note that MixNet does not introduce extra demand monitoring overhead as the state-of-the-art MoE training framework already contains a mechanism to collect this information (meg, [n. d.]) to perform on-demand all-to-all transmission.
@@
5.2. Topology Reconfiguration
In MixNet, finding an optimal topology and deriving an optical schedule is an NP-hard problem (Foerster et al., 2018). We address this challenge by introducing a lightweight greedy algorithm. The key insight is that all-to-all communication time is determined by the delay of largest transfers, which implies that the corresponding GPU pair should be allocated with more circuits. Thus, we identify the communication pairs with the longest transmission time in each iteration and assign them with direct optical links in the OCS topology. The detailed OCS reconfiguration algorithm is shown in Algorithm 1.